# Ambient Diffusion Posterior Sampling: Solving Inverse Problems with Diffusion Models trained on Corrupted Data

### v1

@@ Appendix A Theorem Proof
${\bm{x}}_{t}={\bm{x}}+\sigma_{t}\bm{\eta},\quad{\bm{y}}_{t,\mathrm{train}}=A_{\mathrm{train}}\left({\bm{y}}_{t,\mathrm{train}}+\sigma_{t}\bm{\eta}\right),\quad\tilde{\bm{y}}_{t,\mathrm{train}}=\tilde{A}_{\mathrm{train}}\left(\tilde{\bm{y}}_{t,\mathrm{train}}+\sigma_{t}\bm{\eta}\right).$ (A.3)
@@
We adapt the proof Theorem 4.1 in Ambient Diffusion [17]. To avoid notation clutter, we will be omitting the subscript “train”, when necessary.
@@
From the last equation, it is evident that if $\mathbb{E}[A^{T}A|\tilde{P}]$ is full-rank, and hence the minimizer is $\mathbb{E}[{\bm{x}}_{0}|\tilde{\bm{y}}_{t},\tilde{P}]$ as needed.
Since our $A$ matrix is square, it suffices to show that $\mathbb{E}[A|\tilde{P}]$ is full-rank. By Corollary A.1. in Ambient Diffusion, we have that: $\mathbb{E}[P|\tilde{P}]$ is full-rank. We will now show that $\mathcal{F}^{-1}\mathbb{E}[P|\tilde{P}]\mathcal{F}$ is also full-rank. This can be easily proved with contradiction. Assume that $\mathcal{F}^{-1}\mathbb{E}[P|\tilde{P}]\mathcal{F}$ is not full-rank. Then, there exists a vector ${\bm{w}}\neq 0$ such that:
@@
Appendix B Additional Performance Results
Here we include additional performance results.
@@ Appendix B Additional Performance Results
Appendix C Additional Speed Results
We further include additional speed results.
@@ Appendix C Additional Speed Results
Appendix D Additional MRI Results
Figure 19: Prior samples from Ambient Diffusion trained with under sampled data at $R=2,4,6,8$ (columns $1-4$), EDM trained with L1-wavelet reconstructions of subsampled data at $R=2,4,6,8$ (columns $5-8$), NCSNV2 trained with fully sampled data (column $9$) and EDM trained with fully sampled data (column $10$).
Figure 20: Ambient diffusion one step performance metrics at $R=2,4,6,8$ for models trained at $R=2,4,6,8$.
Figure 21: Ambient diffusion posterior sampling multi-step performance metrics at $R=2,4,6,8$ for models trained at $R=2,4,6,8$.
Figure 22: Performance metrics at $R=2,4,6,8$ for L1-wavelet compressed sensing, diffusion/score-based models trained with fully sampled data (FS-DPS, ALPS), diffusion models trained with L1 reconstructions of subsampled data (L1-DPS)
Figure 24: SSDU MoDL performance metrics at $R=2,4,6,8$ for models trained at $R=2,4,6,8$.
Appendix E Training Hyperparameters for MRI models
The EDM model trained on the fully-sampled ($R=1$) dataset, as well as the four EDM models trained on L1-Wavelet reconstructions at each acceleration factor $R\in\{2,4,6,8\}$ were all trained with 65 million parameters. The Ambient Diffusion models on the other hand were trained with 36 million parameters, for faster training. All the Ambient Diffusion models were trained for 250 epochs. For the Ambient Diffusion training experiments, the further corrupted sampling masks are given as an input to the model by concatenating with the measurements along the channel dimension, as in the original Ambient Diffusion [17] paper.

### v2

@@ A.1 Theory Proof
${\bm{x}}_{t}={\bm{x}}+\sigma_{t}\bm{\eta},\quad{\bm{y}}_{t,\mathrm{train}}=A_{\mathrm{train}}\left({\bm{x}}+\sigma_{t}\bm{\eta}\right),\quad\tilde{\bm{y}}_{t,\mathrm{train}}=\tilde{A}_{\mathrm{train}}\left({\bm{x}}+\sigma_{t}\bm{\eta}\right).$ (A.3)
@@
We adapt the proof Theorem 4.1 in Ambient Diffusion (Daras et al., 2023b). To avoid notation clutter, we will be omitting the subscript “train”, when necessary.
@@
From the last equation, it is evident that if $\mathbb{E}[A^{T}A|\tilde{P}]$ is full-rank, and hence the minimizer is $\mathbb{E}[{\bm{x}}_{0}|\tilde{\bm{y}}_{t},\tilde{P}]$ as needed. Since our $A$ matrix is square, it suffices to show that $\mathbb{E}[A|\tilde{P}]$ is full-rank. By Corollary A.1. in Ambient Diffusion, we have that: $\mathbb{E}[P|\tilde{P}]$ is full-rank. We will now show that $\mathcal{F}^{-1}\mathbb{E}[P|\tilde{P}]\mathcal{F}$ is also full-rank. This can be easily proved with contradiction. Assume that $\mathcal{F}^{-1}\mathbb{E}[P|\tilde{P}]\mathcal{F}$ is not full-rank. Then, there exists a vector ${\bm{w}}\neq 0$ such that:
@@
Note on alternative Ambient Diffusion designs.
Our Ambient Diffusion MRI training approach can be extended to use other linear operators to aggregate the measurements from different coils (e.g. we could have used the pseudoinverse). Alternatively, we can also further condition on the sensitivity maps, or directly feed all the coil measurements as input to the network without aggregation. We opted for a simple and fast method (aggregation through the adjoint). However, these other approaches can lead to improved performance because they fully leverage coil information.
Appendix B Dataset Preparation
We follow standard practices in the MRI literature. We download the FastMRI dataset, randomly select 2,000 T2-Weighted scans, and pick 5 center slices from each scan, giving us us a dataset of 10,000 T2-weighted brain scans ($k$-space measurements). We pass each multi-channel $k$-space sample through a noise pre-whitening filter to transform the noise into standard white Gaussian. Then, we normalize each $k$-space sample by the absolute value of $99$-th percentile of the root sum-of-squares reconstruction of the autocalibration region ($24x\times 24$ pixels from the center). Given the fully sampled $k$-space samples, we estimate sensitivity maps using the ESPIRiT calibration method (Uecker et al., 2014). These measurements serve as a fully sampled reference dataset (i.e., $R=1$).
Appendix C Details for baselines.
L1 Methods.
We used the BART implementation (Uecker et al., 2016; Blumenthal et al., 2022) and searched for the best regularization parameter over the training set. Specifically, we used a regularization weighting of 0.001 with 100 iterations.
Self-Supervised learning via Data Under-sampling (SSDU) (Yaman et al., 2020)
In a supervised setting, end-to-end image restoration networks can be trained by passing measurements through a network like MoDL (Aggarwal et al., 2019) or VarNet (Hammernik et al., 2018) and taking a loss directly on the output of the network with the ground truth image. In the self-supervised setting, however, this is not possible. SSDU, trains the end-to-end restoration network by splitting the available measurement set ($\Omega$) of each sample $y^{\Omega}_{i}$ into training sets ($\Theta$) and loss sets ($\Lambda$) or measurements $y^{\Theta}_{i}$ and $y^{\Lambda}_{i}$ respectively where $\Omega=\Theta\cup\Lambda$. Where the reconstruction network is given access to measurements in the training set to obtain an estimated image $\hat{{\bm{x}}}^{\Theta}_{i}={\bm{h}}_{\theta}(y^{\Theta}_{i})$ and the loss is then defined over the loss measurements as
$L({\bm{y}}^{\Lambda}_{i},\hat{{\bm{x}}}^{\Theta}_{i})=\frac{\left\|{\bm{y}}^{\Lambda}_{i}-A_{\Lambda}\hat{{\bm{x}}}^{\Theta}_{i}\right\|_{1}}{\left\|{\bm{y}}^{\Lambda}_{i}\right\|_{1}}+\frac{\left\|{\bm{y}}^{\Lambda}_{i}-A_{\Lambda}\hat{{\bm{x}}}^{\Theta}_{i}\right\|_{2}}{\left\|{\bm{y}}^{\Lambda}_{i}\right\|_{2}}$ (C.1)
where $A_{\Lambda}$ is the forward operator with ones in the selection mask $P$ at only the measurement locations in the loss set and zeros elsewhere. In this work, we selected ${\bm{h}}_{\theta}(\cdot)$ to have the MoDL architecture (Aggarwal et al., 2019). We trained individual models at four different acceleration levels ($R=2,4,6,8$). Each model was trained for ten epochs using the same training set ($10,000$ slices) as previously described. We used uniform random sampling to separate the training and loss measurement groups for each sample and found that a split of $\rho=\frac{|\Lambda|}{|\Omega|}=0.2$ provided the best performance and thus was used for reporting all subsequent metrics.
MoDL Supervised (MoDL-Sup)
To provide an upper bound on performance for supervised end-to-end methods we also trained the MoDL architecture in a supervised fashion (Aggarwal et al., 2019). Again, we trained individual models at four different acceleration levels ($R=2,4,6,8$). Each model was trained using the same $10,000$ slices as above. However, for these models, we used the normalized root mean squared error (NRMSE) as the loss function:
$L({\bm{x}}_{i},\hat{{\bm{x}}}_{i})=\frac{\left\|{\bm{x}}_{i}-\hat{{\bm{x}}}_{i}\right\|_{2}}{\left\|{\bm{x}}_{i}\right\|_{2}}$ (C.2)
Appendix D Hyperparameters for MRI models
Training Hyperparameters.
The EDM model trained on the fully-sampled ($R=1$) dataset, as well as the four EDM models trained on L1-Wavelet reconstructions at each acceleration factor $R\in\{2,4,6,8\}$ were all trained with 65 million parameters. The Ambient Diffusion models on the other hand were trained with 36 million parameters, for faster training. All the Ambient Diffusion models were trained for 250 epochs. For the Ambient Diffusion training experiments, the further corrupted sampling masks are given as an input to the model by concatenating with the measurements along the channel dimension, as in the original Ambient Diffusion (Daras et al., 2023b) paper.
Sampling Hyperparameters.
The only tunable parameters for DPS (Eq. 2.3) and Ambient DPS (Eq. 2.11) are in the scheduling of the magnitude of the measurements likelihood term. In all the experiments in the DPS paper, this term is kept constant throughout the diffusion sampling trajectory and the authors recommend selecting a value in the range between $[0.1,10]$. We follow this recommendation and we keep this term constant. The value of the step size for each model is selected with a hyperparameter search in the recommended range. For all our experiments, we follow exactly the DPS implementation provided in the official code repository of the paper. The other parameter that impacts performance is the number of steps we are going to run each algorithm for, i.e. the discretization level of the SDEs of Equations 2.3, 2.11. Typically, the higher the number of steps the better the performance since the discretization error decreases (Chen et al., 2022b; 2023). For the performance results, we run each method for several steps $\in\{50,100,150,200,250,300\}$.
Appendix E Additional MRI Results
Table 3: Accelerated MRI reconstruction performance evaluation using LPIPS $\downarrow$ and DISTS $\downarrow$, averaged across 100 test samples.
Training Training Reconstruction $R=2$ $R=4$ $R=6$ $R=8$
Method Data Method LPIPS DISTS LPIPS DISTS LPIPS DISTS LPIPS DISTS
Supervised $R=1$ FS-DPS 0.061 0.028 0.106 0.059 0.166 0.094 0.231 0.126
Self-Supervised $R=2$ A-DPS 0.118 0.062 0.127 0.068 0.138 0.076 0.150 0.082
A-ALD 0.119 0.063 0.127 0.068 0.139 0.077 0.151 0.083
L1-DPS 0.067 0.032 0.131 0.075 0.201 0.115 0.262 0.143
1D-Partitioned SSDU 0.141 0.084 0.144 0.081 0.187 0.108 0.223 0.129
$R=4$ A-DPS 0.155 0.090 0.155 0.092 0.161 0.095 0.168 0.099
A-ALD 0.156 0.090 0.155 0.092 0.161 0.095 0.168 0.099
L1-DPS 0.096 0.051 0.193 0.112 0.258 0.147 0.301 0.167
1D-Partitioned SSDU 0.169 0.105 0.169 0.092 0.187 0.104 0.214 0.123
$R=6$ A-DPS 0.180 0.110 0.178 0.110 0.181 0.111 0.186 0.114
A-ALD 0.179 0.109 0.179 0.111 0.181 0.111 0.186 0.114
L1-DPS 0.112 0.061 0.221 0.128 0.282 0.161 0.318 0.179
1D-Partitioned SSDU 0.143 0.092 0.169 0.102 0.196 0.117 0.222 0.133
$R=8$ A-DPS 0.212 0.129 0.204 0.127 0.204 0.127 0.204 0.127
A-ALD 0.213 0.130 0.204 0.127 0.204 0.127 0.204 0.127
L1-DPS 0.114 0.064 0.227 0.134 0.291 0.168 0.327 0.187
1D-Partitioned SSDU 0.142 0.094 0.176 0.109 0.206 0.127 0.231 0.140
Figure 7: Training samples (left) and prior samples (right) from ambient diffusion model trained on undersampled abdominal MRI scans, where no ground truth data is available.
Figure 8: Ambient diffusion posterior sampling (A-DPS) performance using ambient diffusion model trained on undersampled abdominal MRI scans, where no ground truth data is available.
Figure 9: Ambient diffusion posterior sampling (A-DPS) multi-step performance metrics at $R=2,4,6,8$ for models trained at $R=2,4,6,8$.
Figure 10: Ambient diffusion one step (A-OS) performance metrics at $R=2,4,6,8$ for models trained at $R=2,4,6,8$.
Figure 11: Performance metrics for FS-DPS, L1-DPS, FS-ALD, and L1-CS at $R=2,4,6,8$
Figure 13: SSDU performance metrics at $R=2,4,6,8$ for models trained at $R=2,4,6,8$.
Appendix F Additional Natural Image Performance Results
Figure 14: Speed performance plots for AFHQ.
@@ Appendix F Additional Natural Image Performance Results
For the following experiments, we compare a model trained without corruption to a model trained with randomly inpainted data at $p=0.6$. The evaluation dataset is FFHQ. The training size and hyperparameters for both models are the same. For both models, we extensively tune the sampling hyperparameters to maximize their performance (starting from the recommended hyperparameters from the DPS work). The finding we get in these additional experiments is that the model trained on corrupted data (inpainting) outperforms the model trained on clean data in the high corruption regime across all evaluation tasks.
Figure 19: Box Inpainting results for FFHQ
Figure 20: Additive Gaussian Noise results for FFHQ
Figure 21: Super Resolution + Additive Gaussian Noise ($\sigma=0.05$) results for FFHQ
Appendix G Additional Natural Image Speed Results
@@
Appendix H Limitations and Future Directions.
Our work has several limitations. First, there is a lack of theoretical understanding of the observed experimental phenomenon: we do not have a good grasp of why models trained on highly corrupted data lead to better priors for inverse problems of another corruption type. Second, we only tested the reconstruction performance of our models using the DPS algorithm. Several other recent algorithms have been developed for solving inverse problems with diffusion models and it is unknown whether our findings generalize for these reconstruction algorithms. Finally, our method relies on the existence of Ambient Diffusion models. There are many more models available trained on clean data than models trained on highly corrupted data.
The auto-calibration signal (ACS) region is shared across training samples, which could lead to an over-representation of certain k-space regions. While our current approach does not explicitly address this, introducing a diagonal weighting matrix in the loss function, as suggested by Millard & Chiew (2023), is a promising direction to mitigate this issue.
Currently, we define $A_{train}$ according to Eq 2.7. This is convenient because it lets us use an image-to-image network architecture such as that used by EDM. An alternative could be to define $A_{train}$ directly as in Eq 2.5, in which case the network would go from multi-coil k-space to an image. This network could be preceded by an IFFT, which would effectively be a multi-coil-to-image network, and could be standardized using coil compression (Zhang et al., 2013). Such a network architecture could be interesting and potentially more expressive, as we currently collapse the multi-coil information through zero-filling adjoint. Another possibility is to define $A_{train}$ as the pseudo-inverse applied to $A$, which could be solved efficiently with the conjugate gradient algorithm.
Appendix I Acknowledgements
This research has been supported by NSF Grants CCF 1763702, NSF CCF-2239687 (CAREER), AF1901292, CNS 2148141, Tripods CCF 1934932, NSF AI Institute for Foundations of Machine Learning (IFML) 2019844, the Texas Advanced Computing Center (TACC) and research gifts by Western Digital, WNCG IAP, UT Austin Machine Learning Lab (MLL), Cisco and the Archie Straiton Endowed Faculty Fellowship. Giannis Daras has been supported by the Onassis Fellowship (Scholarship ID: F ZS 012-1/2022-2023), the Bodossaki Fellowship and the Leventis Fellowship.